11-065 Ion Gun Controller Emission Switch Operation

The emission scale switch on the 11-065 ion gun control is used to change the scale of the emission to three levels: 100% (X1), 10%(X.1) and 1%(X.01).  The effect of changing the emission switch is to reduce the emission current, which in turn reduces the ionization (pressure reading) and ion (target) current proportionally.

Initially, the emission needs to be set to 25mA in the X 1 scale and the leak valve is adjusted to achieve 15 to 25 mpA of pressure.

25 mA of emission current
Adjust Argon to get 15 t0 25 mPa of gas pressure

 The graph below shows the target current vs. time and the effect of changing the emission current from X 1 to the X.1 and X.01 scales.

Ion current in uA

When changing the emission scale switch, the emission reading on the 11-065 front panel meter will not change, but the actual emission current will be reduced by a factor of 10 (X.1 scale ) or 100 (X .01 scale). So for example if the emission is set to 25 and the emission scale switch is changed to X .1, the meter still indicates 25mA but the actual emission is 25mA X .1 = 2.5mA.   The pressure display will be reduced by a factor of 10 as well since the number of ions being generated are reduced proportionally to the emission current.  Ideally, the target current will also be reduced by to 10% of the X1 value when the emission scale switch is changed to X.1.  In the graph above we see that the target current drops from about 4.25uA to .8uA as the emission drops from 25mA to 2.5mA.  As expected, the target current drops a factor of 10ish to 70 nA as the emission scale switch is changed to the X .01 position.    

The picture below shows that the pressure drops to about 2mPa as the emission scale is changed to X.1

Emission scale x .1

This picture shows the pressure dropping to .2mPa as the emission scale switch is changed to the X.01 position.    

Emission scale x .01

The emission scale switch is an easy way to reduce the ion target current by a factor of 10 or 100 in order to achieve lower sputter rates.

Note that Balzers RVG 050 thermovalve controllers cannot be used on older 11-065s  in the X .1 or X .01 emission scale positions since they depend on the pressure reading for feedback.  Newer 11-065s have a circuit that compensates so that the pressure output is correct at each emission scale setting.

11-065 High Voltage Arcing problem and solution

As the 11-065s get older we are starting to see instances where the beam voltage, condenser voltage or objective voltage becomes unstable as the front panel potentiometers are adjusted.

The front panel potentiometers can become “noisy” as a result of oxidation on the internal contacts.  In addition to causing instability in the output voltages, it is also possible for the potentiometer output potentiometer “open” up.  When this happens, the output of the beam, condenser or objective supply can go as high as 6.4kV as shown in the picture below in which a 1000:1 high voltage probe is used to measure the beam voltage supply.

In the case of the condenser or objective supplies becoming unstable the result is that the ion beam might go in and out of focus, or the ion beam can get completely shut off.  However, if the beam voltage becomes unstable and goes up to 6.5kV then the opto-couplers on the HV1 board will become damaged and then the emission or pressure circuits will no longer function properly.  The opto-couplers are only rated up to 5kV.

If you suspect that your 11-065 beam, condenser or objective supplies are unstable, here is how to test the HV1 board outputs;

  1. Turn off the 11-065 and unplug the input power cord.
  2. Place the 11-065 on the bench and remove the top cover
  3. Remove the HV cover (on the right hand side of the unit)
  4. Unplug all of the spade connectors which connect the various wires to the HV1 and HV2 boardsHV1 outputs
  5. Lift out the HV2 board (the one closer to the center of the unit)
  6. Place all of the wires off to the side of the 11-065, making sure that none of them touch the chassis. Most of the wires are outputs and so have no voltage on them, but the 4 center wires on the HV2 board have 20VAC on them.
  7. Connect a high voltage probe to the beam output connector on the still plugged in HV1 board, (the board closest to the chassis) ground reference is the chassis. The outputs are:   E93 Beam Voltage, E90 OBJ, E89 COND.Beam voltage HV1 board
  8. Plug in the 11-065 input power cord.
  9. Make sure that the Beam voltage switch is OFF and the beam voltage knob is turned fully CCW.
  10. Turn on the 11-065 power.
  11. Turn on the beam voltage and monitor the output on the DVM that is connected to the high voltage probe.
  12. Slowly turn the Beam voltage potentiometer CW and observe the DVM reading. The Beam voltage output should increase smoothly from near zero to 5kV as you turn up the potentiometer.  If you see jumping, instability, or if the Beam voltage output goes up to 6.5 kV then the potentiometer is noisy and needs to be replaced.  The potentiometer is a 10 k ohm 5 turn 2 Watt 1% potentiometer available from DIgikey, Mouser and Newark.  
  13. The OBJ and COND outputs go from 50% to 100% of the Beam Voltage.  So to test those, the Beam voltage needs to stay fully CW at 5.0   The OBJ and COND potentiometers are also 10 k ohm 5 turn 2 watt 1% potentiometers. 

If the Beam voltage potentiometer was noisy and the voltage went higher than 5kV, then the opto-couplers were likely damaged.   If your do not get any emission current, then most likely U6 on the HV2 board was damaged and should be replaced.  Other components may be damaged as well. If the emission works but not the pressure, U7 is likely damaged.

On the really old 11-065s, U6 was a TIL109 opto.  When that part became obsolete about 20 years ago, it was replaced with a SPX314 opto.  Most of the 11-065s in the field have been updated to the SPX314 (a modification is required).   If you have a really old 11-065 with the TIL109 opto-couplers, RBD still has a few of those in stock.

Recently, the SPX314 has become obsolete and is hard to find.  It can be replaced with a OPI 110 opto-coupler which is also has 15kV of isolation, so even if the HV1 board goes up to 6.4kV this opto-coupler will not get damaged. The OPI 110 will not only repair the problem, it will ensure that this particular problem does not happen again.

There is no modification required to switch from a SPX314 to an OPI 110, but the pin outs are slightly different and are shown in the pictures below.

SPX314 on HV2 board
11-065 opto
OPI 110 opto connections
OPI 110 connection info
U6 schematic HV2 board

The OPI 110 opto-couplers are available from most major electronic part suppliers including Mouser.

Here is a link to the datasheet – OPI 110 datasheet

If you need technical help to repair your 11-065 ion gun control you can contact RBD Instruments Inc. for assistance. We offer technical support, repair/ calibration and the use of a loaner 11-065 while your unit is being repaired.

11-065 Ion Gun Control no high voltage

Over the years I have seen this problem a few times and thought that it would be worth mentioning.

The symptom is that there is no high voltage on the 11-065 ion gun control HV1 board, or the voltage is low.

11-065 Ion Gun Control

11-065 Ion Gun Control

On the system, the symptom will  be no or low target current.

If you have this problem the first thing to check is whether or not C5 on the HV1 board is blown out.   It is an electrolytic capacitor and these days most electrolytic have some creases in them which allow the capacitor to expand and release gas and fluid when it fails.

HV 1 board in 11-065

HV 1 board in 11-065

The location of C5 on the HV1 board is shown below.

C5 location on the HV1 board in an 11-065

C5 location on the HV1 board in an 11-065

For a recent 11-065 repair the symptoms  were no high voltage on the beam and the COND was low. C5  on the HV1 board was obviously blown out so I replaced it.   However I still did not have any high voltage on the beam, and the COND voltage was also low.

The resistors, capacitors, diodes and transorbs on the HV1 board all checked out fine.  It was then that I remembered I have seen this problem once or twice before.   What happened is that when the C5 capacitor failed some of the electrolytic capacitor fluid seeped onto the board.  It was not really noticeable, but there was enough electrolytic capacitor fluid on the board to effectively add a high resistance to the board surface and load down the high voltage supplies.

The schematic for C5 in the filament circuit is shown below.

HV1 board schematic C5

HV1 board schematic C5

The solution was to clean the board with a diluted mixture of Alconox detergent and a tooth brush then rinse the board thoroughly in warm water.  Avoid wetting the transformers.  Blow the HV1 board dry with compressed air and then use a heat gun on low to completely dry the board.

After cleaning all of the electrolytic fluid off and drying the board everything worked fine.

Keep this in mind with any electrolytic capacitor failure where the capacitor blows out and  leaks on the board.

Sign up for RBD Instruments’ Technical Notifications